If Sally can paint a house in 4 hours, and John can paint the same house in 6 hour, how long will it take for both of them to paint the house together?
Answer by Alan3354(30993)

If Sally can paint a house in 4 hours, and John can paint the same house in 6 hour, how long will it take for both of them to paint the house together?
------------
S does 1/4 house per hour
J does 1/6 house per hour
Together, they do 1/4 + 1/6 per hour
= 3/12 + 2/12 = 5/12 house per hour
--> 12/5 hours per house
= 2.4 hours
---------------
Use the shortcut for these, product/sum
4*6/(4+6) = 24/10
= 2.4